# Aristotle’s Second-Order Logic and Natural Deduction

### v17

@@ Abstract
This paper has two goals. The first goal is to show how an extension of second-order logic is a natural framework to formalize portions of Aristotle’s Topics and to bring to the foreground the logical, linguistic and philosophical interest of this work. Aristotelian logic and its related traditions in antiquity are often held to have been equivalent to monadic predicate logic and as such inadequate to formalize mathematics as well as scientific and philosophical discourse in general. The second goal of this paper is to argue that on the contrary the logical theories of Aristotle (which we argue correspond to a variant of natural deduction) and ancient authors such as Galen and Boethius were in fact quite sufficient to account for the logically complex expressions and reasoning involving multiple generality fundamental to the aforementioned disciplines.
@@ 4.4 Additional remarks on Book I of the Topics
Chapter 10. This chapter offers us the rudiments of a new kind of intensional logic: a doxastic logic, and is of considerable interest. We can think of a modal operator Dox(P) satisfying certain logical rules.
@@ Conclusion
Regarding section 6 we can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary epistemic, semantic and logical analysis of our concepts and language - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves232323as we saw in section 6, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’.. The theory of the Categories is already assumed in the Topics. We can ask if Aristotle envisioned the project of a systematic and complete classification of all terms (the entire semantic universe) according to the theory of definition discussed in this paper, starting from the supreme genera. Such a project (which contrasts in certain aspects with that of Leibniz’ characteristica) bears some similarity to the goal of Roget’s famous Thesaurus or of modern lexicological projects.
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\gamma_{2}RS$ (124b)242424if $R$ is one-to-one.

### v18

@@ Abstract
This paper has two goals. The first goal is to show how an extension of second-order logic is a natural framework to formalize portions of Aristotle’s Topics and to bring to the foreground the logical, linguistic and philosophical interest of this work, showing in particular that we are in the presence of a richly intensional and modal conception of logic. Aristotelian logic and its related traditions in antiquity are often held to have been equivalent to monadic predicate logic and as such inadequate to formalize mathematics as well as scientific and philosophical discourse in general. The second goal of this paper is to argue that on the contrary the logical theories of Aristotle (which we argue correspond to a variant of natural deduction) and ancient authors such as Galen and Boethius were in fact quite sufficient to account for the logically complex expressions and reasoning involving multiple generality fundamental to the aforementioned disciplines.
@@ 4.4 Additional remarks on Book I of the Topics
Chapter 10. This chapter offers us the rudiments of a new kind of intensional logic: a doxastic logic, and is of considerable interest. Although we can think of ways of making sense of a doxastic de re modality, it is perhaps better to consider first a de dicto second-order relation $\xi AB$ to be interpreted as: $B$ is generally held to be predicated of $A$ - it is plausible that $B$ is predicatedd of $A$.
@@ Conclusion
ESOL (in its standard form or with bounded quantifiers232323and Dependent Type Theory, also called Constructive Type Theory.) does not in itself formalize extensional logic. It need not even come with equality predicates (there must be more than one: equality must be divided between a binary first-order predicate for individual equality and second-order binary predicates for first-order $n$-ary predicates). Even allowing a standard Leibnizian axiomatics of equality, Frege still had to introduce extensionalism as something derived from an ad hoc axiom (Law V). It follows from this law that two predicates are equal iff they apply to the same elements (they have the same ’extension’ or ’course-of-values’).
In the post-Fregean history of philosophical logic there were schools which posited extensionalism as a dogma and strove furthermore to ban intensionality and modality from logic altogether or to explain away these in an extensionalist anner. Other logicians, mathematicians and philosophers (including Husserl himself) viewed extensionalism and Frege’s Law V as the ultimate source of Russell’s paradox and a main source of other paradoxes and confusions in the philosophy of logic (this is explored in particular in [8]). We agree entirely with the thesis of [12] that one can view the remarkable developments in 20th century logic (and I would include homotopy type theory) as a progressive liberation from the shackles of extensionalism and (revidindicating Husserl’s program) a gaining of awareness of the essential, inescapable, intensional and modal aspects of logical systems including classical logic itself (further evidenced by the theory of control operators and the $\lambda\mu$ -calculus). I would add that the current view would be that extensionalism does not flow from the nature of logic itself but is something added and posited for the needs of certain specific ’regional ontologies’ . A major thesis of our paper is that in Aristotle we find what amounts to a rich purely intensional extended second-order logic and at the same time a more specialized development of an extensionalist logic (including with a possible worlds semantics) in the Prior Analytics. Thus both Frege and post-Fregean developments are likewise anticipated.
A major aspect of the present work is the thesis that Aristotle’s theory of logic in the Topics is much more general, richer and philosophically insightful than the subsequent more specialized syllogistic theory of the Prior Analytics with the possible exception of some technical aspects of the modal syllogistic. For monadic first-order predicates, our arguments amount to showing that a major difference lies in that second-order relations used in the Prior Analytics are extensional while those used in the Topics are more fine-grained and intensional. The transition between the second-order relations of the Topics (such as $\gamma$, $\iota$) and that of the Analytics is one of a flattening and collapse wherein the modal differentiations, the different ways of a predicate belonging to another predicate, are discarded in favor of a purely extensional criterion of predication ($\kappa$). But the extended second-order logic of the Topics is essentially an intensional logic.
Having an intensional extended second-order logic allows us to formalize a potentially unlimited variety of de dicto modalities by using second-order relations $\rho AB$ applied to predicates $A$ and $B$, expressing how $B$ is related to $A$ independently of the extensions of these predicates. De re modalities can likewise be formalized through second-order functions as briefly discussed in Section 3.2. In this discussion we point out how this alternative approach can also interpret our AML system we developed for the modal syllogistic (which however used a variant of standard extensionalist possible worlds semantics). It may be that the a non-extensional extended second-order approach to de re modalities for monadic predicates could help refine this interpretation.
Returning to de dicto modalities, we explored a little in Section 5 the richness of these modes present in Aristotle’s Topics. Aristotle defines a second-order quaternary relation which cannot be interpreted except as a precursor of probabilistic logic. We wrote the relation as $\mathbb{P}_{\iota}AB>\mathbb{P}_{\iota}A^{\prime}B^{\prime}$ but it is understood that this is a single quaternary relation $\mathbb{P}ABA^{\prime}B^{\prime}$ which expresses that $B$ is more likely to be the property of $A$ than $B^{\prime}$ is to be the property of $A^{\prime}$. Aristotle then presents topoi which express several interresting axioms concerning this relation.
Also (as we saw when we discussed the problemsf of the identity of a term and its definition or the treatment of identity in Book VII of the Topics) a major avenue of exploration concerns the different fine-grained (in particular qualitative) concepts of equality and identity (which include the very particular and specialized notion of extensional equality).
In our investigations of the complex textual and interpretative problems surrounding the relations $\delta,\Delta$ and accident $\sigma$ (specially in Section 4.1) the variety of kinds of intensional modes of predication becomes apparent as well as the connection to the standard modalities (i.e. those that appear in the modal syllogistic and which are already slightly richer than the standard modern ones). Although we cannot offer yet a definite account, it appears to be a kind of ontological gradation starting from the center and progressing to the periphery of a being or concept. At the heart is the relation of definition expressing core essence, then follows essential and inseparable accidents, then property (which does not express essence and yet is inseparable) and finally different kinds of separable ’accident’ which are defined by contingency, potentiality and genericity (something being the case most of the time). There is also a relation of having a predicate ’naturally’ or ’by nature’. It would be very interesting to consider these relations in light of Husserl’s mereology proposed in the Third Logical Investigation.
@@
Regarding Section 6 we can of course question if this dialectical game could lead to actual progress in philosophy or to certain or plausible knowledge about reality. In particular the exact scope of $K$ seems to be vague and uncertain; the assumption of the possession of a perfectly disambiguated language is itself implausible (although it should be said that the Topics is acutely aware of this problem). However we should not go to the other extreme of discarding this kind of philosophical practice. Rather what was required is a preliminary epistemic, semantic and logical analysis of our concepts and language - and such concerns are in no way foreign to the mature Aristotle: things better known to us are not necessarily better known in themselves242424as we saw in Section 6, the extra-logic second-order relation of ’being better known’ plays an important role in the Topics: if $A$ is a property of $B$ then it has to be better known than $B$. In book I,1 Aristotle says that a syllogism is a ’proof’ when the premises are ’true and primitive’ or epistemically depending on ’true and primitive’ premises. The adjective ’alethê’ is also translated as ’evident’. Aristotle then clarifies that evident and primitive is what is known through itself and not through another’. All ’scientific principles’ should be self-evident. Other propositions are ’endoksa’, translated sometimes as ’probable’.. The theory of the Categories is already assumed in the Topics. We can ask if Aristotle envisioned the project of a systematic and complete classification of all terms (the entire semantic universe) according to the theory of definition discussed in this paper, starting from the supreme genera. Such a project (which contrasts in certain aspects with that of Leibniz’ characteristica) bears some similarity to the goal of Roget’s famous Thesaurus or of modern lexicological projects.
@@
9. $\sim\forall xSx(\iota yR^{\ast}xy)\rightarrow\sim\gamma_{2}RS$ (124b)252525if $R$ is one-to-one.